OpenStudy (anonymous):

solve abs(x+2)<=8+abs(2x-4)

OpenStudy (anonymous):

 so always

OpenStudy (anonymous):

statement is true for all x in the real numbers

OpenStudy (anonymous):

how about solve analytically, please

OpenStudy (jdoe0001):

hmmm so let's check the vertices you have 2 absolute values, so you have 2 vertices what makes |x+2| to 0? well x = -2, one vertex what makes |2x-4| to 0? well x = 2, 2nd vertex so we end up with 3 intervals to values against 

OpenStudy (anonymous):

where's the 8?

OpenStudy (jdoe0001):

\(\bf \textit{left side -, right side -}\\ -x-2 \le 8-2x+4\\ \textit{setting x = -3}\\ 3-2 \le 8+6+4 \textit{ \# true, thus is an answer}\)

OpenStudy (jdoe0001):

ohh, the 8 is a shift, but in the graph is not used, just the vertices is all I used and the graph the expression itself

OpenStudy (anonymous):

i see what you're doing...

OpenStudy (jdoe0001):

and doing the test for the the 2nd interval and 3rd interval with - and + and + and + for the expressions, will yield all of them are true, thus are solutions